Each exterior angle of the regular n-gon has a measure of 5 degrees. find n
Answer by ikleyn(52775)

.
Use the fact that the sum of all interior angles of any (convex) polygon is equal to 360 degrees.


It gives you an equation


    n*5 = 360,


with the solution


    n = 360/5 = 72.      ANSWER